OpenStudy (anonymous):

Will input question once this opens with the Equation Button was allowed at the beginning. Will input question once this opens with the Equation Button was allowed at the beginning. @Mathematics

OpenStudy (anonymous):

Solve each inequality. Express your answer in interval notation. (Please show me the work in this I am still confused on how to get the answers.) Thanks for any assistance. \[3x^2+8x \le0\]

OpenStudy (mathmagician):

\[x(3x+8)=0\] therefore x=0 and x=-8/3. When x<(-8/3) or x>0 \[3x^2+8x>0\] therefore the interval is [-8/3;0]

OpenStudy (anonymous):

How do you know which direction the inequality goes?

OpenStudy (anonymous):

thought it would be x>-8/3 and x<0

OpenStudy (anonymous):

It should be x < -8/3 or x > 0

OpenStudy (mathmagician):

take any value from interval [-8/3;0] and you will see that 3x^2+8x<=0 and if you take for example 2, 3x^2+8x will be >0

OpenStudy (anonymous):

that's not the way it should be derived ..

OpenStudy (mathmagician):

but still x>0 or x<-8/3 is not truth

OpenStudy (anonymous):

I haven't read the whole thing .. I meant that \[ 3x^2+8x>0 \Rightarrow \left.x<-\frac{8}{3}\right\|x>0 \]

OpenStudy (mathmagician):

then it would be right :)

OpenStudy (anonymous):

and for the first thing \[ -\frac{8}{3}\leq x\leq 0 \]

OpenStudy (anonymous):

The problem started as \[3x^2+8x \le0\]

OpenStudy (anonymous):

btw why you did it like that .. this is so simple .. just factor the main thing and you will see you answer instantly!

OpenStudy (mathmagician):

if i did for myself, i'd have factored, but the way i did is better for explanation

OpenStudy (anonymous):

How so? :)

OpenStudy (anonymous):

I factored it but couldn't figure out which direction the > or < symbol went. I am guessing it has to do with the division switching the direction of the > <.

OpenStudy (anonymous):

No .. say for < 0 two one can positive and one have to be negative ..